# UAV-Mounted IRS (UMI) in the Presence of Hovering Fluctuations: 3D Pattern Characterization and Performance Analysis

### v2

@@ II System Model
where depend on variables $\varepsilon_{x}$ and $\varepsilon_{y}$. Considering the influence of all IRS elements, the normalized radiation pattern of the IRS towards angles $\tilde{\theta}_{t}$ and $\tilde{\phi}_{t}$ from IRS to BS, along with $\tilde{\theta}_{r}$ and $\tilde{\phi}_{r}$ from IRS to UE is defined as [28]
$\mathbb{G}(\tilde{\theta}_{t},\tilde{\phi}_{t},\tilde{\theta}_{r},\tilde{\phi}_{r})=\mathbb{G}_{e}(\tilde{\theta}_{t},\tilde{\phi}_{t},\tilde{\theta}_{r},\tilde{\phi}_{r})\times\mathbb{G}_{a}(\tilde{\theta}_{t},\tilde{\phi}_{t},\tilde{\theta}_{r},\tilde{\phi}_{r}),$ (3)
@@
where $\mathbb{G}_{a}(.)$ is the normalized array factor and $\mathbb{G}_{e}(.)$ is the single element radiation pattern. Because of UAV fluctuations, it is deduced from (3) that the normalized radiation pattern of IRS depends on two independent RVs $\varepsilon_{x}$ and $\varepsilon_{y}$. The single element radiation pattern is obtained as [28]
$\mathbb{G}_{e}(\tilde{\theta}_{t},\tilde{\phi}_{t},\tilde{\theta}_{r},\tilde{\phi}_{r})=E(\tilde{\theta}_{t},\tilde{\phi}_{t})\times E(\tilde{\theta}_{r},\tilde{\phi}_{r}),$ (4)
where $E(\theta,\phi)$ is calculated as
$E(\theta,\phi)=\begin{cases}\cos^{3}(\theta),&\theta\in[0,\pi/2],\phi\in[0,2\pi]\\ 0,&\theta\in(\pi/2,\pi],\phi\in[0,2\pi].\end{cases}$ (5)
@@
Furthermore, the normalized array factor can be expressed as (6) (See top of next page) where $\delta_{1}=-\sin({{\theta}}_{t})\cos({{\phi}}_{t})-\sin({{\theta}}_{r})\cos({{\phi}}_{r})$ and $\delta_{2}=-\sin({{\theta}}_{t})\sin({{\phi}}_{t})-\sin({{\theta}}_{r})\sin({{\phi}}_{r})$ [28] .
@@
Note that when the UAV experiences no fluctuations, $\mathbb{G}_{a}=1$, whereas during fluctuations, $\mathbb{G}_{a}<1$, leading to a decrease in the power of the reflected signal directed to UE.
II-B Channel Model
All channels between IRS and BS/UE are modeled as Rician distributed, with Rician factor $K_{i}=A^{2}/(2\delta^{2})$ where $i=0,1$. Here, $K_{0}$ and $K_{1}$ correspond to the BS-IRS and IRS-UE links, respectively. The scale parameter is defined as $B=A^{2}+2\delta^{2}$, where $A^{2}$ and $2\delta^{2}$ represent the power of the line-of-sight (LoS) and non-line-of-sight (NLoS) components. Assuming a normalized average received power, $B$ is set to $1$. The PDF of a Rician random variable is given as follows [29]
$f_{Z}(z)=\frac{z}{\delta^{2}}\exp(-\frac{z^{2}+A^{2}}{2\delta^{2}})I_{0}(\frac{zA}{\delta^{2}}),z>0,$ (7)
where $I_{0}(.)$ represents the modified Bessel function of the first kind with order zero.
II-C General Received Signal Model
In Fig. 1, the channels between BS and IRS, as well as between IRS and UE are denoted by $\bar{\mathbf{H}}_{BI}$ and $\bar{\mathbf{h}}_{IU}$, respectively, which can be modeled as [30],
$\bar{\mathbf{H}}_{BI}=\sqrt{\beta_{0}}\mathbf{H}_{BI},$ (8)
$\bar{\mathbf{h}}_{IU}=\sqrt{\beta_{1}}\mathbf{h}_{IU}.$
where $\beta_{0}$ and $\beta_{1}$ represent path loss components, and $\mathbf{H}_{BI}\in\mathbb{C}^{N\times M}$ and $\mathbf{h}_{IU}\in\mathbb{C}^{N\times 1}$ denote small-scale fading of links BS-IRS and IRS-UE, respectively. It is assumed that the small-scale fading remains constant within each coherence interval. Additionally, perfect CSI is assumed for all IRS-related links. Accordingly, $\beta_{0}$ and $\beta_{1}$ are defined as:
$\beta_{k}\triangleq c_{0}{d_{k}}^{-\alpha_{k}},\quad k\in\{0,1\},$ (9)
where $c_{0}$ represents the reference path loss at a distance of $1$m. The path loss exponents for BS-IRS and IRS-UE links are denoted by $\alpha_{0}$ and $\alpha_{1}$, respectively, whereas $d_{0}$ and $d_{1}$ represent the distance from the IRS to the BS and UE, respectively. The reflection matrix of the active IRS is defined as $\mathbf{\Theta}\triangleq\mathbf{A}\mathbf{\Phi}$, where $\mathbf{\Phi}\triangleq\text{diag}(e^{j\phi_{1}},...,e^{j\phi_{N}})$ represents the IRS phase shift matrix with $\phi_{n}$ being the phase shift at the $n$-th element, $n\in\{1,2,...,N\}$. Additionally, the active IRS amplification matrix is denoted by $\mathbf{A}\triangleq\text{diag}(a_{1},...,a_{N})$ where $a_{n}$ is the amplification factor of the $n$-th element. Unlike passive IRS, active IRS introduces non-negligible thermal noise at each reflector, denoted as $\mathbf{n}_{F}\in\mathbb{C}^{N\times 1}$, which is typically modeled as $\mathbf{n}_{F}\sim\mathcal{N}(\mathbf{0}_{N},N_{F}\mathbf{I}_{N})$, where $N_{F}$ is the amplification noise power [10]. Finally, the received signal at UE is expressed as
$y=\sqrt{\mathbb{G}}{\bar{\mathbf{h}}}^{H}_{IU}\mathbf{\Theta}{\bar{\mathbf{H}}}_{BI}\mathbf{w}s+\sqrt{\mathbb{G}}{\bar{\mathbf{h}}}^{H}_{IU}\mathbf{\Theta}\mathbf{n}_{F}+n,$ (10)
where $n\sim\mathcal{N}(0,N_{0})$ is the additive white Gaussian noise (AWGN) at the UE with power $N_{0}$. The transmit signal from BS is represented by $s$ with transmit power $P_{t}$. The beamforming vector is denoted by $\mathbf{w}\in\mathbb{C}^{M\times 1}$. The power constraint at active IRS is expressed as follows [31]
$P_{t}\|\mathbf{A\Phi}{\bar{\mathbf{H}}}_{BI}\mathbf{w}\|^{2}+\|\mathbf{A\Phi}\mathbf{I}_{N}\|^{2}N_{F}\leq P_{F},$ (11)
where $P_{F}$ is the maximum amplification power of the active IRS. Therefore, the instantaneous SNR at UE in the general model, i.e., MISO-Active-IRS, can be expressed as
$\text{SNR}=\frac{P_{t}\mathbb{G}|{\bar{\mathbf{h}}}^{H}_{IU}\mathbf{A\Phi}{\bar{\mathbf{H}}}_{BI}\mathbf{w}|^{2}}{\mathbb{G}\|\bar{\mathbf{h}}^{H}_{IU}\mathbf{A\Phi}\|^{2}N_{F}+N_{0}}.$ (12)
@@ III IRS Radiation pattern Distribution Function
The SNR in (12) depends on the IRS radiation pattern, $\mathbb{G}$. To analysis the performance, it is necessary to have the distribution of $\mathbb{G}$. Therefore, in this section, we derive the PDF of the IRS radiation pattern. From (3), it follows that $\mathbb{G}=\mathbb{G}_{e}\times\mathbb{G}_{a}$. To obtain the distribution of $\mathbb{G}$, we perform three steps. First, we derive the distribution of $\mathbb{G}_{a}$, and subsequently, we derive the distribution of $\mathbb{G}_{e}$. Finally, we obtain the distribution of $\mathbb{G}$ by applying the results from the last two steps.
@@
$Z_{x}\triangleq\sin({\tilde{\theta}}_{t})\cos({\tilde{\phi}}_{t})+\sin({\tilde{\theta}}_{r})\cos({\tilde{\phi}}_{r})+\delta_{1},$ (13)
$Z_{y}\triangleq\sin({\tilde{\theta}}_{t})\sin({\tilde{\phi}}_{t})+\sin({\tilde{\theta}}_{r})\sin({\tilde{\phi}}_{r})+\delta_{2},$
@@
where $g_{ax}$ and $g_{ay}$ are the array factors along the $x$-axis and $y$-axis, respectively. For a function of one variables $f(x)$ whose first-order derivative $f_{x}$ exists at a point $a$, the $1^{\footnotesize\text{st}}$-order Taylor polynomial near the point $a$ is
@@
We define $\varepsilon_{\theta_{q}}(\varepsilon_{x},\varepsilon_{y})\triangleq{\tilde{\theta}}_{t}(\varepsilon_{x},\varepsilon_{y})-\theta_{t}$ and $\varepsilon_{\phi_{q}}(\varepsilon_{x},\varepsilon_{y})\triangleq{\tilde{\phi}}_{t}(\varepsilon_{x},\varepsilon_{y})-\phi_{t}$, representing the fluctuations of the elevation and azimuth angles, respectively. Utilizing (15), we can approximate the values of $\sin(\tilde{x})$ and $\cos(\tilde{x})$ in (13) with $\tilde{x}=x+\varepsilon$, where $\varepsilon$ is near zero, as follows
$\sin(\tilde{x})\approx\sin(x)+\varepsilon\cos(x),$ (16)
$\cos(\tilde{x})\approx\cos(x)-\varepsilon\sin(x).$
@@
$Z_{x}(\varepsilon_{\theta_{t}},\varepsilon_{\phi_{t}},\varepsilon_{\theta_{r}},\varepsilon_{\phi_{r}})\approx\cos\theta_{t}\cos\phi_{t}\varepsilon_{\theta_{t}}$ $-\sin\theta_{t}\sin\phi_{t}\varepsilon_{\phi_{t}}$ (17)
$+\cos\theta_{r}\cos\phi_{r}\varepsilon_{\theta_{r}}$ $-\sin\theta_{r}\sin\phi_{r}\varepsilon_{\phi_{r}},$
@@
$+\cos\theta_{r}\sin\phi_{r}\varepsilon_{\theta_{r}}$ $+\sin\theta_{r}\cos\phi_{r}\varepsilon_{\phi_{r}}.$
@@
$\mu_{\varepsilon_{\theta_{q}}}=A_{\theta_{qx}}\mu_{x}+A_{\theta_{qy}}\mu_{y},$ (18)
$\sigma^{2}_{\varepsilon_{\theta_{q}}}=A^{2}_{\theta_{qx}}\sigma^{2}_{x}+A^{2}_{\theta_{qy}}\sigma^{2}_{y},$
@@
$\mu_{\varepsilon_{\phi_{q}}}=A_{\phi_{qx}}\mu_{x}+A_{\phi_{qy}}\mu_{y},$ (20)
$\sigma^{2}_{\varepsilon_{\phi_{q}}}=A^{2}_{\phi_{qx}}\sigma^{2}_{x}+A^{2}_{\phi_{qy}}\sigma^{2}_{y},$
$A_{\phi_{qx}}=-\frac{(1+\tan^{2}\theta_{qx})\tan\theta_{qy}}{\tan^{2}\theta_{qx}+\tan^{2}\theta_{qy}},$ (21)
$A_{\phi_{qy}}=-\frac{(1+\tan^{2}\theta_{qy})\tan\theta_{qx}}{\tan^{2}\theta_{qx}+\tan^{2}\theta_{qy}}.$
@@
Since all RVs in (17), i.e., $\varepsilon_{\theta_{t}}$, $\varepsilon_{\phi_{t}}$, $\varepsilon_{\theta_{r}}$ and $\varepsilon_{\phi_{r}}$, follow Gaussian distributions individually, and any linear combination of them also follows a Gaussian distribution, it can be concluded that they are jointly Gaussian RVs. To illustrate, consider the linear combination of them for arbitrary coefficients $\alpha$, $\beta$, $\zeta$ and $\kappa$ using (22) and (90) (See Appendix A) as follows
$\alpha\varepsilon_{\theta_{t}}$ $+\beta\varepsilon_{\phi_{t}}+\zeta\varepsilon_{\theta_{r}}+\kappa\varepsilon_{\phi_{r}}=\alpha A_{\theta_{tx}}\varepsilon_{x}+\alpha A_{\theta_{ty}}\varepsilon_{y}$ (23)
$+\beta A_{\phi_{tx}}\varepsilon_{x}+\beta A_{\phi_{ty}}\varepsilon_{y}+\zeta A_{\theta_{rx}}\varepsilon_{x}+\zeta A_{\theta_{ry}}\varepsilon_{y}$
@@
As clear from (23), this linear combination is itself a linear function of the independent Gaussian RVs $\varepsilon_{x}$ and $\varepsilon_{y}$. Therefore, any such linear combination remains Gaussian. Therefore, the distributions of $Z_{k}$ for $k\in\{x,y\}$ follows the Gaussian distribution, specifically $Z_{k}\sim\mathcal{N}(\mu_{Z_{k}},\sigma^{2}_{Z_{k}})$ where $\mu_{Z_{k}}$ is calculated as
@@
Following a similar approach, the approximation of $g_{ay}$ can be derived in the same manner as (28) by replacing the subscript $x$ with $y$. Now, we split $g_{ax}$ in (28) into several sectors and propose a simpler model given by
$g_{ax}\simeq\Pi(\frac{D_{a}N_{q}|Z_{x}|}{2})+\sum_{i=1}^{lD_{a}-1}{\frac{D_{a}^{2}(1-\cos(\frac{2\pi i}{D_{a}}))}{2\pi^{2}i^{2}}}$ (29)
×$\qquad\qquad\times[\Pi(\frac{D_{a}N_{q}|Z_{x}|}{2(i+1)})-\Pi(\frac{D_{a}N_{q}|Z_{x}|}{2i})],$
where $\Pi(x)=\begin{cases}1,&\mbox{if }x\leq 1\\ 0,&\mbox{if }x>1\end{cases}$. $D_{a}$ represents the number of sectors, and $l\in\{1,2\}$ whereby $l=1$ corresponds considering only the main lobe of the pattern, while $l=2$ is used for higher precision by including both the main-lobe and the first side-lobe. The simplified model for $g_{ay}$ can be obtained similarly to (29) by replacing the subscript $x$ with $y$. Fig. 3 plots the exact model and sectoral model of the IRS array factor along $x$-axis, $g_{ax}$ versus $\varepsilon_{x}$ for $D_{a}=5$ and $15$, with $N=64$. We can increase the accuracy of the sectoral model by increasing $D_{a}$, however, this improvement comes at the cost of increased computational complexity. As shown in Fig. 3, when $l=1$, the proposed model only takes into account the main lobe, whereas, for $l=2$, the first side lobe is also included.
@@
$f_{\mathbb{G}_{a}}(\mathbb{G}_{a})=\sum_{i=1}^{J_{a}}p_{a}(i)\delta(\mathbb{G}_{a}-q_{a}(i))$ (30)
where $J_{a}=l^{2}D_{a}^{2}$, $\mathbf{q}_{a}=\mathbf{q}_{ax}\otimes\mathbf{q}_{ay}$ and $\mathbf{p}_{a}=\mathbf{p}_{ax}\otimes\mathbf{p}_{ay}$. Here, $\mathbf{q}_{ak}\in\mathbb{R}^{lD_{a}\times 1}$ and $\mathbf{p}_{ak}\in\mathbb{R}^{lD_{a}\times 1}$ where
$q_{ak}(i)=\frac{D_{a}^{2}(1-\cos(\frac{2\pi i}{D_{a}}))}{2\pi^{2}i^{2}},$ (31)
$p_{ak}(i)=Q(\frac{2i-D_{a}N_{q}\mu_{Z_{k}}}{D_{a}N_{q}\sigma_{Z_{k}}})-Q(\frac{2(i+1)+D_{a}N_{q}\mu_{Z_{k}}}{D_{a}N_{q}\sigma_{Z_{k}}})$ (32)
$\qquad+Q(\frac{2i+D_{a}N_{q}\mu_{Z_{k}}}{D_{a}N_{q}\sigma_{Z_{k}}})-Q(\frac{2(i+1)-D_{a}N_{q}\mu_{Z_{k}}}{D_{a}N_{q}\sigma_{Z_{k}}})$ ,
$k\in\{x,y\}$
@@
Step2: Deriving the distribution of $\mathbb{G}_{e}$
@@
where $g_{et}$ and $g_{er}$ are the single element radiation patterns in direction of angles $\theta_{t}$ and $\theta_{r}$. Then, by dividing $g_{eq}$ in (33) where $q\in\{t,r\}$ with multiple sectors, we propose a simple model given by
$g_{eq}\simeq\cos^{3}$ $(\theta_{q})\Pi(LD_{e}|\varepsilon_{\theta_{q}}|)+\sum_{i=(1-D_{e})/2}^{(D_{e}-1)/2}{\cos^{3}(\theta_{q}+\frac{2i}{LD_{e}})}$ (34)
×$\qquad\quad\times\text{sgn}(i)[\Pi(\frac{LD_{e}\varepsilon_{\theta_{q}}}{2i+1})-\Pi(\frac{LD_{e}\varepsilon_{\theta_{q}}}{2i-1})],$
where
$\text{sgn}(x)=\begin{cases}-1,&\mbox{if }x<0\\ 0,&\mbox{if }x=0\\ 1,&\mbox{if }x>0\end{cases}.$ (35)
$D_{e}$ is the number of sectors and $L$ is utilized to adjust the considered range of variation of $\varepsilon_{x}$ and $\varepsilon_{y}$. As $L$ increases, a smaller range of variation for $\varepsilon_{x}$ and $\varepsilon_{y}$ is considered. Additionally, as $L$ increases, a consistent number of sectors results in more accuracy. Fig. 4 plots the sectoral model and exact model of IRS single radiation pattern versus $\varepsilon_{x}$ for $D_{e}=5$ and $9$. Increasing $D_{e}$ increases the accuracy of the proposed model at the cost of more complexity. Consequently, selecting the optimal value for $D_{e}$ requires balancing acceptable computational complexity with the desired level of accuracy. Moreover, increasing $L$ reduces the range of variations that need to be considered, resulting in fewer sectors that can be accurately reached.
Figure 4: The exact model of $g_{eq}$, given in (33) and its sectoral model, proposed in (34) versus $\varepsilon_{x}$.
Lemma 4.
The PDF of single element radiation pattern $G_{e}$ is derived as:
$f_{\mathbb{G}_{e}}(\mathbb{G}_{e})=\sum_{i=1}^{J_{e}}p_{e}(i)\delta(\mathbb{G}_{e}-q_{e}(i))$ (36)
where $J_{e}=D_{e}^{2}$, $\mathbf{q}_{e}=\mathbf{q}_{et}\otimes\mathbf{q}_{er}$ and $\mathbf{p}_{e}=\mathbf{p}_{et}\otimes\mathbf{p}_{er}$. Here, $\mathbf{q}_{ek}\in\mathbb{R}^{D_{e}\times 1}$ and $\mathbf{p}_{ek}\in\mathbb{R}^{D_{e}\times 1}$ where
$q_{ek}(i)=\cos^{3}(\theta_{k}+\frac{2i}{LD_{e}}),$ (37)
$p_{ek}(i)=Q(\frac{2i-1-LD\mu_{\varepsilon_{\theta_{k}}}}{LD\sigma_{\varepsilon_{\theta_{k}}}})-Q(\frac{2i+1-LD\mu_{\varepsilon_{\theta_{k}}}}{LD\sigma_{\varepsilon_{\theta_{k}}}})$ , (38)
$k\in\{t,r\}$
Proof.
Please refer to Appendix C. ∎

### v3

@@ II-B Received Signal Model
where $\mathbb{G}$ describes the array radiation pattern of the IRS towards angles $\tilde{\theta}_{t}$ and $\tilde{\phi}_{t}$ from IRS to BS, along with $\tilde{\theta}_{r}$ and $\tilde{\phi}_{r}$ from IRS to UE, scaling the received power at the user, defined as [41]
$\mathbb{G}(N_{q},\zeta_{t},\zeta_{r})=\mathbb{G}_{e}(\zeta_{t},\zeta_{r})\times\mathbb{G}_{a}(N_{q},\zeta_{t},\zeta_{r}),$ (16)
@@
where $\zeta_{t}=\{\tilde{\theta}_{t},\tilde{\phi}_{t}\}$, $\zeta_{r}=\{\tilde{\theta}_{r},\tilde{\phi}_{r}\}$. Here, $\mathbb{G}_{a}(.)$ is the normalized array factor and $\mathbb{G}_{e}(.)$ is the single-element radiation pattern. Because of UAV fluctuations, it is deduced from (16) that the normalized radiation pattern of IRS depends on two independent RVs $\varepsilon_{x}$ and $\varepsilon_{y}$. The single-element radiation pattern is obtained as [41]
$\mathbb{G}_{e}(\zeta_{t},\zeta_{r})=E(\zeta_{t})\times E(\zeta_{r}),$ (17)
@@
where $E(\zeta)$ for $\zeta=\{\theta,\phi\}$ is calculated from (6). Furthermore, the normalized array factor can be expressed as (18) shown at the top of next page, where $\delta_{1}=-\sin({{\theta}}_{t})\cos({{\phi}}_{t})-\sin({{\theta}}_{r})\cos({{\phi}}_{r})$ and $\delta_{2}=-\sin({{\theta}}_{t})\sin({{\phi}}_{t})-\sin({{\theta}}_{r})\sin({{\phi}}_{r})$ [41].
@@
Remark 1.
When the UAV is stable (i.e., experiences no fluctuations), $\mathbb{G}_{a}=1$, whereas during fluctuations (e.g., due to wind or dynamic movement), $\mathbb{G}_{a}<1$, leading to a decrease in the power of the reflected signal directed to the UE. Furthermore, as the number of IRS elements, $N$, increases, the IRS pattern becomes narrower. While this enhances directivity under stable conditions, it also makes the system more sensitive to UAV fluctuations. Even minor UAV fluctuations can cause significant beam misalignment, significantly reducing $\mathbb{G}_{a}$. Thus, a fundamental trade-off exists between employing larger IRS arrays to achieve higher gain and maintaining robustness against UAV fluctuations.
Remark 2.
The instantaneous SNR for a passive IRS is obtained by $A^{*}=1$ and $\sigma_{f}^{2}=0$. Substituting these values into (15) yields
$\gamma_{\text{max}}=\frac{P_{t}M(1-\zeta)\beta_{0}\beta_{1}\mathbb{G}(\sum_{n=1}^{N}|H_{BI,m,n}||h_{IU,n}|)^{2}}{P_{t}\zeta\sigma_{e}^{2}+\sigma_{n}^{2}}.$ (19)
@@ III IRS Radiation Pattern Distribution Function
The maximum SNR in (15) depends on the IRS radiation pattern, $\mathbb{G}$. To analyze the performance, it is necessary to have the distribution of $\mathbb{G}$. Therefore, in this section, we derive the PDF of the IRS radiation pattern. From (16), it follows that $\mathbb{G}=\mathbb{G}_{e}\times\mathbb{G}_{a}$. To obtain the distribution of $\mathbb{G}$, we perform three steps. First, we derive the distribution of $\mathbb{G}_{a}$, and subsequently, we demonstrate that UAV fluctuations induce negligible changes in $\mathbb{G}_{e}$, thus it can be treated as deterministic. Finally, we obtain the distribution of $\mathbb{G}$ by applying the results from the last two steps.
@@
$Z_{x}\triangleq\sin({\tilde{\theta}}_{t})\cos({\tilde{\phi}}_{t})+\sin({\tilde{\theta}}_{r})\cos({\tilde{\phi}}_{r})+\delta_{1},$ (20a)
$Z_{y}\triangleq\sin({\tilde{\theta}}_{t})\sin({\tilde{\phi}}_{t})+\sin({\tilde{\theta}}_{r})\sin({\tilde{\phi}}_{r})+\delta_{2},$ (20b)
@@
where $g_{ax}$ and $g_{ay}$ are the array factors along the $x$-axis and $y$-axis, respectively. To approximate $Z_{x}$ and $Z_{y}$ in (21) defined in (20), we employ linear approximation. For a function of one variable $f(x)$ whose first-order derivative $f_{x}$ exists at a point $a$, the 11nd-order Taylor polynomial near the point $a$ is
@@
[…]
$\sin(\tilde{x})\approx\sin(x)+\varepsilon\cos(x),$ (23a)
$\cos(\tilde{x})\approx\cos(x)-\varepsilon\sin(x).$ (23b)
@@
$Z_{x}(\varepsilon_{\theta_{t}},\varepsilon_{\phi_{t}},\varepsilon_{\theta_{r}},\varepsilon_{\phi_{r}})$ $\approx\cos\theta_{t}\cos\phi_{t}\varepsilon_{\theta_{t}}-\sin\theta_{t}\sin\phi_{t}\varepsilon_{\phi_{t}}$
$+\cos\theta_{r}\cos\phi_{r}\varepsilon_{\theta_{r}}-\sin\theta_{r}\sin\phi_{r}\varepsilon_{\phi_{r}},$ (24a)
@@
$+\cos\theta_{r}\sin\phi_{r}\varepsilon_{\theta_{r}}+\sin\theta_{r}\cos\phi_{r}\varepsilon_{\phi_{r}}.$ (24b)
Note that the linear approximation, which enabled our analytical derivation, is valid under small-angle variation conditions (1–2 degrees), a state maintained by stable UAV flight controllers. Our simulation results under these conditions confirm the model accuracy, showing only $1\%$ approximation error for $\sigma_{k}^{2}=1^{\circ},k\in\{x,y\}$, which is negligible and does not affect the key insights or conclusions of the paper. A limitation of this approach is that its accuracy may degrade for highly dynamic flight maneuvers with larger attitude variances.
To derive the PDF of $\mathbb{G}_{a}$, we must determine the distribution of $Z_{x}$ and $Z_{y}$, which depend on $\varepsilon_{\theta_{q}}$ and $\varepsilon_{\phi_{q}}$, whose distributions are determined in Lemma 26 and Lemma 28.
@@
$\mu_{\varepsilon_{\theta_{q}}}=A_{\theta_{qx}}\mu_{x}+A_{\theta_{qy}}\mu_{y},$ (25a)
$\sigma^{2}_{\varepsilon_{\theta_{q}}}=A^{2}_{\theta_{qx}}\sigma^{2}_{x}+A^{2}_{\theta_{qy}}\sigma^{2}_{y},$ (25b)
@@
$\mu_{\varepsilon_{\phi_{q}}}=A_{\phi_{qx}}\mu_{x}+A_{\phi_{qy}}\mu_{y},$ (27a)
$\sigma^{2}_{\varepsilon_{\phi_{q}}}=A^{2}_{\phi_{qx}}\sigma^{2}_{x}+A^{2}_{\phi_{qy}}\sigma^{2}_{y},$ (27b)
$A_{\phi_{qx}}=-\frac{(1+\tan^{2}\theta_{qx})\tan\theta_{qy}}{\tan^{2}\theta_{qx}+\tan^{2}\theta_{qy}},$ (28a)
$A_{\phi_{qy}}=-\frac{(1+\tan^{2}\theta_{qy})\tan\theta_{qx}}{\tan^{2}\theta_{qx}+\tan^{2}\theta_{qy}}.$ (28b)
@@
Since all RVs in (24), i.e., $\varepsilon_{\theta_{t}}$, $\varepsilon_{\phi_{t}}$, $\varepsilon_{\theta_{r}}$ and $\varepsilon_{\phi_{r}}$, follow Gaussian distributions individually, and any linear combination of them also follows a Gaussian distribution, it can be concluded that they are jointly Gaussian RVs. To illustrate, consider the linear combination of them for arbitrary coefficients $\alpha$, $\beta$, $\xi$, and $\kappa$ using (29) and (80) (See Appendix A) as follows
$\alpha\varepsilon_{\theta_{t}}$ $+\beta\varepsilon_{\phi_{t}}+\xi\varepsilon_{\theta_{r}}+\kappa\varepsilon_{\phi_{r}}=\alpha A_{\theta_{tx}}\varepsilon_{x}+\alpha A_{\theta_{ty}}\varepsilon_{y}$ (30)
$+\beta A_{\phi_{tx}}\varepsilon_{x}+\beta A_{\phi_{ty}}\varepsilon_{y}+\xi A_{\theta_{rx}}\varepsilon_{x}+\xi A_{\theta_{ry}}\varepsilon_{y}$
@@
It is clear from (30) that this linear combination is itself a linear function of the independent Gaussian RVs $\varepsilon_{x}$ and $\varepsilon_{y}$. Therefore, any such linear combination remains Gaussian. Therefore, the distributions of $Z_{k}$ for $k\in\{x,y\}$ follows the Gaussian distribution, specifically $Z_{k}\sim\mathcal{N}(\mu_{z_{k}},\sigma^{2}_{z_{k}})$ where $\mu_{z_{k}}$ is calculated as
@@
Following a similar approach, the approximation of $g_{ay}$ can be derived in the same manner as (35) by replacing the subscript $x$ with $y$. Deriving a closed-form PDF for $\mathbb{G}_{a}$ is mathematically intractable. We cannot easily integrate over this complex function to find a statistical distribution. To solve this, we propose a sectoral model, which is a form of simplification. Therefore, we split $g_{ax}$ in (35) into several sectors and propose a simpler model given by
$g_{ax}\simeq\Pi(\frac{DN_{q}|Z_{x}|}{2})+\sum_{i=1}^{lD-1}{\frac{D^{2}(1-\cos(\frac{2\pi i}{D}))}{2\pi^{2}i^{2}}}$ (36)
$\qquad\qquad\times[\Pi(\frac{DN_{q}|Z_{x}|}{2(i+1)})-\Pi(\frac{DN_{q}|Z_{x}|}{2i})],$
where $\Pi(x)=\begin{cases}1,&\mbox{if }x\leq 1\\ 0,&\mbox{if }x>1\end{cases}$. $D$ represents the number of sectors, and $l\in\{1,2\}$ whereby $l=1$ corresponds to considering only the main lobe of the pattern, while $l=2$ is used for higher precision by including both the main lobe and the first side lobe. The simplified model for $g_{ay}$ can be obtained similarly to (36) by replacing the subscript $x$ with $y$. Fig. 3 plots the exact model and sectoral model of the IRS array factor along the $x$-axis, $g_{ax}$ versus $\varepsilon_{x}$ for $D=5$ and $15$, with $N=64$. We can increase the accuracy of the sectoral model by increasing $D$; however, this improvement comes at the cost of increased computational complexity, as noted in Remark 6. As shown in Fig. 3, when $l=1$, the proposed model only takes into account the main lobe, whereas, for $l=2$, the first side lobe is also included.
@@
$f_{\mathbb{G}_{a}}(\mathbb{G}_{a})=\sum_{i=1}^{J}p_{a}(i)\delta(\mathbb{G}_{a}-q_{a}(i))$ (37)
where $J=l^{2}D^{2}$, $\mathbf{q}_{a}=\mathbf{q}_{ax}\otimes\mathbf{q}_{ay}$ and $\mathbf{p}_{a}=\mathbf{p}_{ax}\otimes\mathbf{p}_{ay}$. Here, $\mathbf{q}_{ak}\in\mathbb{R}^{lD\times 1}$ and $\mathbf{p}_{ak}\in\mathbb{R}^{lD\times 1}$ where
$q_{ak}(i)=\frac{D^{2}(1-\cos(\frac{2\pi i}{D}))}{2\pi^{2}i^{2}},$ (38)
$p_{ak}(i)=Q(\frac{2i-DN_{q}\mu_{Z_{k}}}{DN_{q}\sigma_{Z_{k}}})-Q(\frac{2(i+1)+DN_{q}\mu_{Z_{k}}}{DN_{q}\sigma_{Z_{k}}})$ (39)
$\qquad+Q(\frac{2i+DN_{q}\mu_{Z_{k}}}{DN_{q}\sigma_{Z_{k}}})-Q(\frac{2(i+1)-DN_{q}\mu_{Z_{k}}}{DN_{q}\sigma_{Z_{k}}})$ ,
$k\in\{x,y\}$ .
@@
Step2: Treating $\mathbb{G}_{e}$ as a Deterministic Component
@@
where $g_{et}$ and $g_{er}$ are the single element radiation patterns in direction of angles $\theta_{t}$ and $\theta_{r}$. For a function of one variable $f(x)$ whose first- and second-order derivatives $f^{(1)}(x)$ and $f^{(2)}(x)$, respectively, exist at the point $a$, the 22nd-order Taylor polynomial near the point $a$ is
$f(x)\approx f(a)+f^{(1)}(a)(x-a)+f^{(2)}(a)(x-a)^{2}.$ (41)
Using (41), for small $\varepsilon_{\theta_{q}}$, the second-order Taylor expansion around $\theta_{q}$ is
$g_{eq}\approx\cos^{3}(\theta_{q})$ $-3\cos^{2}(\theta_{q})\sin(\theta_{q})\varepsilon_{\theta_{q}}$ (42)
$+\frac{3}{2}\cos(\theta_{q})(2-3\cos^{2}(\theta_{q}))\varepsilon^{2}_{\theta_{q}}$
We approximate $\mathbb{G}_{e}\approx\cos^{3}(\theta_{t})\cos^{3}(\theta_{r})$. The expected error of this approximation is
$E\{\text{error}\}=$ $-3\cos^{2}(\theta_{q})\sin(\theta_{q})E\{\varepsilon_{\theta_{q}}\}$ (43)
$+\frac{3}{2}\cos(\theta_{q})(2-3\cos^{2}(\theta_{q}))E\{\varepsilon^{2}_{\theta_{q}}\}$
Given that $E\{\varepsilon_{\theta_{q}}\}=0$ and $E\{\varepsilon^{2}_{\theta_{q}}\}$ is very small, the expected error is negligible. Moreover, $\mathbb{G}_{e}$ is independent of the number of IRS elements $N$, in contrast to the normalized array factor $\mathbb{G}_{a}$, which is highly sensitive to $N$. For the subsequent optimization over $N$, we therefore treat $g_{eq}$ as the constant $g_{eq}\approx\cos^{3}(\theta_{q})$. Specifically, the $\mathbb{G}_{e}$ is treated as
$\mathbb{G}_{e}=\cos^{3}(\theta_{t})\times\cos^{3}(\theta_{r}).$ (44)
This treatment is consequently applied to $E(\zeta_{q})$ for $q\in\{t,r\}$, i.e., $E(\zeta_{q})=\cos^{3}(\theta_{q})$.